what is percntage of sulphur in CN2H4S ?

Grade:12

2 Answers

Gaurav
8 years ago
Molar mass of CN2H4S = 12+28+4+32 = 76
Percentage of Sulphur = (32/76)x 100 = 42%
sajid
8 years ago
%COMPONENT=(WEIGHT OF COMPONENT/MOLAR MASS OF MOLECULE)*100
Weight of compoent(sulphur)=32
Molar mass of molecule(CN2H4S)=76          (i.e., 12+2(14)+4(1)+32=76 )
%sulphur = (32/76)*100 =42% approx.